OpenStudy (anonymous):

Given the complex number, find a cube root of the number. It should be exact and in the form z^1/3 = r(cos(theta) + i sin(theta)) z = 4 + i4*sqrt(3)... HELP!!!

OpenStudy (anonymous):

\[z=4+i4 \sqrt{3}\]

OpenStudy (amistre64):

r = the distance form 0+0i to 4+4sqrt(3) i

OpenStudy (anonymous):

Need in \[z ^{1/3}=r*\cos(\theta)+i \sin(\theta)\]

OpenStudy (anonymous):

Okay...

OpenStudy (amistre64):

\[z^n=r^n(cos(nx)+isin(nx))\]

OpenStudy (amistre64):

r = sqrt(4^2+(4sqrt(3))^2)

OpenStudy (amistre64):

lol or,given z=a+bi\[r=\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

why replace z with r?

OpenStudy (amistre64):

becuase z is a complex number; r is the modulus of z which is analologulous to the distance from the origin of (a,b)

OpenStudy (amistre64):

analagus?

OpenStudy (amistre64):

anomoluos?

OpenStudy (anonymous):

you are above my knowledge now. i am just taking trig.

OpenStudy (amistre64):

this is trig :)

OpenStudy (anonymous):

we haven't covered modulus.

OpenStudy (anonymous):

i know what it means.

OpenStudy (anonymous):

it is remainder.

OpenStudy (anonymous):

but i am trying to rework trig now that i am in group theory, but we have done nothing with complex.

OpenStudy (amistre64):



OpenStudy (anonymous):

so what is z^1/3???

OpenStudy (anonymous):

If you recognize the factor of four times the one and square root of three in the coefficients, you can divide by eight to get\[z = 4 + i4\sqrt(3)=8(1/2+i \sqrt{3}/2)=8(\cos(\pi/3)+isin(\pi/3))\]

OpenStudy (anonymous):

is that r^1/3

OpenStudy (amistre64):

z^(1/3) = r^(1/3) * (cos(x/3) + i sin(x/3))

OpenStudy (amistre64):

i dont think pi enters into the picture there

OpenStudy (amistre64):

\[z=a+bi\] \[a=rcos(x);\ bi=irsin(x)\]

OpenStudy (amistre64):

\[z=rcos(x)+irsin(x)\] \[z=r(cos(x)+isin(x))\] \[z^{1/3}=(r(cos(x)+isin(x)))^{1/3} \] \[z^{1/3}=r^{1/3}\ (cos(x)+isin(x)))^{1/3} \]

OpenStudy (anonymous):

From that, we can get \[z ^{1/3}=2(\cos \theta+isin \theta)\]\[\theta=\pi/9\pm2\pi/3\]

OpenStudy (amistre64):

oh yeah, we got an actual angle dont we :)

OpenStudy (anonymous):

okay. i am just SO confused with this trig. I am retaking all the stuff from the class 2 years ago and trying to make sense of it. Thanks!

OpenStudy (amistre64):

good luck with it :)

OpenStudy (anonymous):

Thanks!